I have worked on this problem trying to figure it out.
When all of the gorcery carts escape from the supermarket, it takes Reginald 12 minutes to round them up and bring them back. Because Norman doesn't make as much per hour as Reginald, it takes norman 18 minutes to do the same job. How long would it take them working together to complete the round up?
I know that Rate x Time = work
I believe that in setting this up, it would have to be Time = Work/Rate. However you would have to know the Work though, right?

Found 2 solutions by stanbon, alka001:
Answer by stanbon(75887)

it takes Reginald 12 minutes to round them up and bring them back. Because Norman doesn't make as much per hour as Reginald, it takes norman 18 minutes to do the same job. How long would it take them working together to complete the round up?
-------------------------------------------
Reginald DATA:
time = 12 min/job ; rate=1/12 job/min
--------------------
Norman DATA:
time = 18 min/job ; rate = 1/18 job/min
----------------------
Working together DATA:
time = x ; rate = 1/x
-----------------------------
EQUATION:
Reg rate + Norman rate = together rate
1/12 + 1/18 = 1/x
30/216=1/x
x= 216/30 = 7.2 minutes (number of minutes for them to do the job together)
-----------------
Cheers,
Stan H.

Answer by alka001(14)

it is a problem on time (unitary method)
Reginald does the job in 12 min.
So, his 1 minutes job would be 1/12
Norman does the job in 18 min.
So his 1 minutes job is 1/18
When they both work together,
their 1 min job would be1/12 +1/18 =(18+12)/12*18 =30/12*18 =5/36
So, if both work together they would do the job in 36/5 minutes =7min 12 sec